# ResQuNNs:Towards Enabling Deep Learning in Quantum Convolution Neural Networks

### v3

@@ Abstract
In this paper, we present a novel framework for enhancing the performance of Quantum Convolution Neural Networks, a.k.a Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and addressing the critical challenges associated with them. Traditional quanvolutional layers, although beneficial for feature extraction, have largely been static, offering limited adaptability. Unlike state-of-the-art, our research overcomes this limitation by enabling training within these layers, significantly increasing the scaling and learning potential of QuNNs. However, in deep QuNNs, i.e., QuNNs with multiple trainable quanvolutional layers induce complexities in gradient-based optimization, primarily due to the restricted gradient flow across these layers. To resolve this, we propose a novel architecture, Residual Quanvolutional Neural Networks (ResQuNNs), leveraging the concept of residual learning, which facilitates the flow of gradients by adding skip connections between these layers. The insertion of residual blocks between quanvolutional layers ensures enhanced gradient access throughout the network, thus enabling deep learning in QuNNs and eventually leading to improved training performance. Moreover, we provide empirical evidence on the strategic placement of these residual blocks within ResQuNNs. Through extensive experimentation, we identify efficient configurations of residual connections, which enable gradients flow across all the layers in the network that eventually results in efficient training. Our results mark a substantial step forward in the evolution of quantum deep learning, offering new avenues for both theoretical development and practical quantum computing applications.
@@ I Introduction
There are two primary variants of QCNNs, each leveraging quantum mechanics principles like superposition, entanglement, and interference. The first type, Quantum-Inspired CNNs, proposed in [3], mirrors the structure of classical CNNs but replaces convolutional and pooling layers with deeper PQCs. This design demands higher qubit coherence to fully exploit quantum parallelism and entanglement, posing a challenge for current quantum hardware. Nevertheless, a significant research has been dedicated to explore this very architectural framework for various applications [5, 10]. The second type, Quanvolutional Neural Networks (QuNNs), first introduced in [11], and visually depicted in Fig. 1, focuses on quantum-enhanced convolutional layers. These networks encode classical image pixels into quantum states, apply unitary transformations via PQCs, and then measure the qubits to create output image channels, which are subsequently processed by classical fully connected layers. This approach offers a more feasible implementation with current quantum technology.
QuNNs have widely been used in the literature for various applications such as 2D and 3D radiological image classification [12], analysing the high energy physics data [13], for object detection and classification [14, 15]. In light of the capabilities and constraints of present-day quantum hardware, this paper also delves into QuNNs. Our primary focus is on identifying the challenges faced in training QuNNs and proposing effective solutions to address these issues.
Figure 1: A typical Illustration of Quanvolutional Neural Network Architecture.
@@
I-A Motivational Analysis
QuNNs, as shown in Fig. 1 performs the convolution on the input image within the grey-shaded region. This process involves sequentially encoding a small region of the input image into quantum space, while moving a kernel across the image, indicated by the green-shaded area of Fig. 1. A unitary transformation is then applied to a small square region of the input via the PQC a.k.a ansatz, as depicted in the brown region of Fig. 1. The output of this quanvolution (a quantum circuit-based convolution) is then processed by a classical neural network to produce the final network output. While this architecture is standard in many state-of-the-art works, a key challenge is the untrainable nature of the quanvolution operation, primarily used for feature extraction. Consequently, the learning process is primarily driven by the subsequent classical network. For QuNNs to be fully scalable, ensuring the trainability of underlying quanvolutional layers is crucial.
In Fig. 2, we assess the impact of trainable versus untrainable quanvolutional layers in QuNNs. To this end, we train QuNNs with both trainable (allowing weight updates during backpropagation) and untrainable (frozen weights, no updates) quanvolutional layers. This training is performed on $1000$ randomly chosen images from MNIST dataset over $30$ iterations, using an $80\%-20\%$ training-validation split. The Adam optimizer with a $0.01$ learning rate was employed. We utilized a $2\times 2$ kernel resulting in $4$-qubit quanvolutional layers, with each layer comprising of $4$ parameterized $R_{y}$ gates and $4$ $CNOT$ entanglement gates. The depth of quantum circuit in the quanvolutional layer is set to $5$.
@@ I-A Motivational Analysis
While a single trainable quanvolutional layer enhances QuNNs’ training efficacy, as depicted in Fig. 2, scalability of QuNNs requires multiple trainable layers. Also, the rationale for using multiple layers is to avoid the problem of vanishing gradients a.k.a Barren Plateaus (BP), common in single, more expressive PQCs with a higher qubit count and sufficient depth[16, 17, 18]. Multiple layers, with progressively smaller quantum circuits, can potentially mitigate the BPs[19]. However, challenges emerge when multiple quanvolutional layers are trained concurrently, as gradients are exclusively accessible only in the last quanvolutional layer, thus limiting the optimization process only to this layer. Fig. 3 demonstrates this by showing gradients for two and three cascaded trainable quanvolutional layers.
We observe that the gradients of only the final quanvolutional layer are accessible, while those of the preceding layers are effectively None. The restricted gradient flow means that in multi-layered QuNNs, only the final layer contributes in the learning process, leaving earlier layers remain uninvolved in optimization. This presents a significant challenge in effectively training multi-layered QuNNs, thereby necessitating innovative approaches to address this challenge in the pursuit of scalable and effective QuNN architectures.
@@ I-B Novel Contributions
  • Trainable Quanvolutional Layers: We introduce the concept of making quanvolutional layers trainable in QuNNs, addressing the challenge of enhancing their adaptability within QuNNs and improving their training performance.
@@
II ResQuNN Methodology
@@
II-A Trainable Quanvolutional Layers in QuNNs.
The first key aspect of our methodology involves the concept of making quanvolutional layers trainable. The majority of works in the literature utilizes these layers for efficient feature extraction and pass them to classical layers, however, the quanvolution operation itself is not trainable. We recognize the need to enhance the adaptability of these layers within QuNNs and introduce trainable quanvolutional layers to harness their full potential. Making the quanvolutional layer trainable entails making the quantum circuit (labeled as 1 in Fig. 5) trainable. We do this by making the quanvolution layer as a pytorch layer which is later used as a standard layer during the model construction. The trainable QuNN utilized in this study is depicted in purple shaded region of Fig. 5, and operates through the following steps.
@@
    where $l$ represents the total number of layers, i.e., number of times the circuit is repeated before measurement. The size of output after every quanvloutional layers is detailed in Table I. |$\ket{\psi_{\text{QCL1}}}$ is the final state state after the first quanvolutionla layer before measurement and upon measurement the quantum state collapses and produce a classical result:
@@
    First, we use classical layer to postprocess the output of quanvolutional layer, as highlighted in label 2 in Fig. 5). This is a major practice in state-of-the-art QuNN architectures, where a fully connected neuron layer is used to take the data from the quanvolutional layer and produces the final classification output. The number of neurons in this layer are equal to the total number of classes in the dataset used, which in our case is $10$ since we used MNIST dataset which has $10$ classes (more details in Section III).
    Let’s assume the output of the quanvolutional layer is a set of measurements $\{M=m_{1},m_{2}\ldots m_{n}\}$, where each $m_{i}$ represents the measurement of a qubit in terms of classical information, such as the expectation values in computational basis. This result is transformed to a vector format and then passed to the fully connected layer with $10$ neurons. The output feature size from quanvolutional layers is different and depends on the kernel size and stride similar to that of classical convolutional layers. We restrict our experimentation to kernel size of $2\times 2$ and stride of $2$, and output feature size in case of different residual configuration (label 5 in Fig. 5) is presented in the Table I.
    TABLE I: Output Feature Size for Two Quanvolutional Layer
@@
    The classical neural network layer with $10$ neurons can be represented as a fully connected layer, where each neuron receives inputs from all $n$ measurements of the quanvolutional layer and processes these inputs to produce an output. The output of $j$-th neuron in the classical layer can be described by the following Equation:
@@
    We also use quantum circuit to postprocess the result from quanvolutional layer, as highlighted by label 3 in Fig. 5. This is primarily for two reasons: firstly to demonstrate the working of fully quantum QuNNs, and secondly, to better understand the importance of residual connections by reducing the influence of classical layers being used at the end.
    The use of a quantum circuit for post-processing introduces certain computational overheads, mainly due to the prerequisite of re-encoding the (classical) outputs from the quanvolutional layers, a process that entails considerable computational expense. Also the following quantum circuit after the encoding takes longer to produce the results than a simple classical neuron layer. We use $4$ classes from MNIST dataset for experimental simplification. For the purpose of encoding the outputs from the quanvolutional layer, Amplitude Encoding [20], is used which can be described by the following Equation:
@@
    where $x_{i}$ is the $i^{th}$ element of $x$ and |$\ket{i}$ is the $i^{th}$ computational basis state. We used amplitude encoding primarily because it allows to encode $2^{n}$ features in $n$ qubits. Based on Table I, the smallest output feature size is $7\times 7\times 1=49$, which requires atleast $49$ qubits via angle encoding (one qubit per feature) and only $6$ qubits via amplitude encoding. Similarly, the maximum output size of quanvolutional layer is $28\times 28\times 1=784$, requiring $784$ qubits. Angle encoding in this case would require $784$ qubits, whereas Amplitude Encoding reduces the requirement to merely $10$ qubits. The use of $784$ qubits is not practical in current NISQ devices, therefore we use amplitude encoding to encode the qunavolutional layer’s result for further processing by the quantum circuit. To ensure a consistent impact of the quantum circuit’s post-processing across all residual configurations, a $10$-qubit circuit is utilized for processing the outputs of the quanvolutional layer, corresponding to the maximal feature size of $784$. However, as mentioned earlier, we use $4$ classes, therefore, only $4$ qubits are measured each corresponding to one class. The encoded data is then subjected to the unitary evolution,which can be described by the following Equation:
@@
II-B Residual Quanvolutional Neural Networks (ResQuNNs).
To overcome the gradient accessibility issue encountered in multi-layered QuNNs, as discussed in Section I-A, we introduce ResQuNNs. Drawing inspiration from classical residual neural networks, ResQuNNs incorporate residual blocks between quanvolutional layers. The residual blocks combine the output of existing quanvolutional layers with their own or the previous layer’s input before forwarding it to the subsequent quanvolutional layer, as shown in label 4 in Fig. 5, and play a pivotal role in facilitating comprehensive gradient access and, in turn, improving the overall training performance of QuNNs. The integration of residual blocks ensures that gradients flow efficiently across the network, thereby overcoming the challenges associated with training multiple quanvolutional layers.
II-C Strategic Placement of Residual Blocks.
Within the ResQuNN architecture, not every random residual configuration results in comprehensive gradient flow throughout the network, and residual blocks needs to be strategically placed between quanvolutional layers. In our empirical investigation, we conduct extensive experimentation to determine the appropriate locations for inserting residual blocks within ResQuNNs comprising two quanvolutional layers, as shown in Fig. 5, labeled 5. This analysis helps us uncover the most effective residual configurations for ResQuNNs, enabling the gradients to flow throughout the proposed ResQuNN architecture, resulting in a deeper architecture and and eventually the enhanced training performance. Based on the results of two quanvolutional layers, only the most effective configurations, i.e., those facilitating gradient flow through both layers were chosen for subsequent examination within three-layer quanvolutional frameworks, as denoted by label 6 in Fig. 5.
We then trained ResQuNNs with two quanvolutional layers for all residual configurations for a multiclass classification problem (more on dataset in Section III) to assess the impact of gradient accessibility, both throughout the entire network and within segments of the network, on the learning efficacy of QuNNs.
TABLE II: Dataset and Hyperparemeters Specifications. BS = batch size, Opt = optimizer and LR = learning rate
@@
III Experimental Setup
We trained the proposed ResQuNNs on a selected subset of the MNIST dataset. Table II summarizes the dataset and hyperparameters specification used in this paper. The ResQuNNs are trained for $30$ epochs. Adam optimizer with learning rate of $0.01$ is used for the optimization. The objective function chosen for training is cross entropy loss, which can be described (for C classes) by the following equation:

### v4

@@ Abstract
In this paper, we introduce a framework aimed at enhancing the performance of Quantum Convolutional Neural Networks a.k.a Quanvolutional Neural Networks (QuNNs) by incorporating trainable quanvolutional layers and addressing the significant trainability challenges inherent in multi-layered QuNNs. Current state-of-the-art QuNNs predominantly utilize static quanvolutional layers, which limit their feature extraction capabilities. Our proposed approach mitigates this limitation by enabling the training of these layers, thereby substantially improving the scalability and learning potential of QuNNs. However, deep QuNNs—those with multiple trainable quanvolutional layers—encounter complexities in gradient-based optimization, primarily due to restricted gradient flow across these layers. To resolve this, we propose Residual Quanvolutional Neural Networks (ResQuNNs), leveraging the concept of residual learning, which facilitates the flow of gradients by adding skip connections between these layers. The insertion of residual blocks between quanvolutional layers ensures enhanced gradient access through all the layers of the network, thus enabling deep learning in QuNNs and eventually leading to an improved training performance. Moreover, we provide empirical evidence on the strategic placement of these residual blocks within ResQuNNs. Through extensive experimentation, we identify efficient configurations of residual connections, which enable gradients flow across all the layers in the network that eventually results in efficient training. Our results mark a substantial step forward in the evolution of quantum deep learning, offering new avenues for both theoretical development and practical quantum computing applications.
Index Terms:
Quantum neural networks, Residual learning, Trainability, deep learning
@@ I Introduction
There are two primary variants of QCNNs, each leveraging quantum mechanics principles like superposition, entanglement, and interference. The first type, Quantum-Inspired CNNs, proposed in [3], mirrors the structure of classical CNNs but replaces convolutional and pooling layers with deeper PQCs. This design demands higher qubit coherence to fully exploit quantum parallelism and entanglement, posing a challenge for current quantum hardware. Nevertheless, a significant research has been dedicated to explore this very architectural framework for various applications [5, 10]. The second type, Quanvolutional Neural Networks (QuNNs), first introduced in [11], and visually depicted in Fig. 1(a), focuses on quantum-enhanced convolutional layers. These networks encode classical image pixels into quantum states, apply unitary transformations via PQCs, and then measure the qubits to create output image channels, which are subsequently processed by classical fully connected layers. This approach offers a more feasible implementation with current quantum technology. Therefore, in this paper, we also delve into QuNNs with a focus on identifying the challenges faced in training QuNNs and proposing effective solutions to address these issues.
Figure 1: (a) A typical Illustration of Quanvolutional Neural Network Architecture, (b) Simplified Workflow of Quanvolutional Layer in QuNNs
@@
An abstract view of the workflow of quanvolution layer in QuNNs is presented in Fig. 1(b), which is four step process: (1) the original image $U$ is divided into spatially located subsections $u_{x}$. Each subsection is a $d\times d$ matrix with $d>1$, (2) these image subsections are then encoded into a quantum state ($q$), e.g., for an encoding function $E$, the quantum state can be defined as $q_{x}=E(u_{x})$. We encode the input features into the qubit rotation angles using $R_{y}(\theta)$. (3) the encoded state $q_{x}$ then undergoes a unitary transformation via a quantum circuit a.k.a ansatz $Q$ which can be described as; $f_{x}=Q(q_{x})=Q(E(u_{x}))$, (4) decoding is then performed via measurements which can be described as, $D(f_{x})=D(Q(E(u_{x})))$. By repeating these steps for different subsections, the full input image is traversed, producing an outcome organized as a multi-channel image. The output of this quanvolution is usually processed by a classical neural network to produce the final output.
This QuNN architecture, widely adopted in state-of-the-art research, is extensively utilized for various applications such as 2D and 3D radiological image classification [12], high energy physics data analysis [13], and object detection and classification [14, 15]. In these contexts, the quantum circuit in QuNNs primarily functions as a feature extractor and is not trained during the quanvolution operation. Consequently, the learning process is primarily driven by the subsequent classical network.
Although there are recent studies that employ trainable quantum circuits in quanvolutional layers, to the best of our knowledge, no existing literature addresses multi-layered QuNNs. Table I some summarizes recent works utilizing QuNNs with both trainable and non-trainable quanvolutional layers. To achieve full scalability in QuNNs, it is crucial to ensure the trainability of underlying quanvolutional layers in multi-layered architectures.
TABLE I: Recent State-of-the-art QuNNs employing trainable and untrainable quanvolution layers
Ref Trainable Quanvolutional Layers Multiple Quanvolutional Layers
[16] $\times$ $\times$
[17] $\times$ $\times$
[18] $\times$ $\times$
[19] $\times$ $\times$
[20] $\times$ $\times$
[21] $\times$ $\times$
[22] $\checkmark$ $\times$
[23] $\checkmark$ $\times$
I-A Motivational Case Study
In Fig. 2, we assess the impact of trainable versus untrainable quanvolutional layers in QuNNs. To this end, we train QuNNs with both trainable (allowing weight updates during backpropagation) and untrainable (frozen weights, no updates) quanvolutional layers. This training is performed on $1000$ randomly chosen images from MNIST dataset over $30$ iterations, using an $80\%-20\%$ training-validation split. The Adam optimizer with a $0.01$ learning rate was employed. We utilized a $2\times 2$ kernel resulting in $4$-qubit quanvolutional layers, with each layer comprising of $4$ parameterized $R_{y}$ gates and $3$ $CNOT$ gates for (nearest neighbor) qubit entanglement. The overall depth of quantum circuit in the quanvolutional layer is set to $2$, accounting to a total of $8$ single and $6$ two-qubit gates.
@@ I-A Motivational Case Study
While a single trainable quanvolutional layer enhances QuNNs’ training efficacy, as depicted in Fig. 2, scalability of QuNNs requires multiple trainable layers. However, challenges emerge when multiple quanvolutional layers are trained concurrently, as gradients are exclusively accessible only in the last quanvolutional layer, thus limiting the optimization process only to this layer. Fig. 3 demonstrates this by showing gradients for two and three trainable quanvolutional layers.
We observe that the gradients of only the final quanvolutional layer are accessible, while those of the preceding layers are effectively None. The restricted gradient flow means that in multi-layered QuNNs, only the last quanvolutional layer contributes in the learning process, leaving earlier layers remain uninvolved in optimization. This presents a significant challenge in effectively training multi-layered QuNNs, thereby necessitating innovative approaches to address this challenge for scalable and effective QuNN architectures.
@@ I-B Novel Contributions
  • Trainable Quanvolutional Layers: We first make the quanvolutional layers trainable in QuNNs, addressing the challenge of enhancing their adaptability within QuNNs and improving their training performance.
@@
  • Both Classical and Quantum Postprocessing: State-of-the-art QuNNs utilize classical networks to postprocess the output of quanvolutional layers. However, to demonstrate the effectiveness of the residual approach in multi-layered QuNNs (with no influence on the overall learning from classical layers), other than classical postprocessing layer, we employ the quantum circuit itself for postprocessing the results of the quanvolutional layers.
II Related Work
Recently, some studies have explored the application of residual connections in quantum neural networks (QNNs). In [24], residual connections are incorporated into a simple feedforward QNN architecture to address the problem of barren plateaus (BPs). BPs are characterized by the exponential vanishing of gradients in parameterized quantum circuits, which are essential components of QNNs, as the number of qubits increases. By partitioning the conventional QNN architectures into multiple nodes, and then by adding residual connections between these nodes, [24] successfully mitigates the BP issue. Our work diverges from [24] in two significant ways: (1) We employ residual approach in a different architecture, specifically QuNNs, and (2) we address a distinct underlying problem, i.e., restricted gradient flow in QuNNs, which occurs regardless of the number of qubits unlike BPs which occurs with increasing number of qubits. To overcome the partial gradient flow in QuNNs, we apply residual connections.
In another recent work [25], the concept of residual connections is used in QNNs. The proposed quantum residual neural networks (QResNets) leverage auxiliary qubits in data-encoding and trainable blocks to extend frequency generation forms and improve the flexibility of adjusting Fourier coefficients, leading to enhanced spectral richness and expressivity of parameterized quantum circuits at the cost of greater qubit utilization in the form of auxiliary qubits, which is not suitable for NISQ devices. In contrast, our work focuses on QuNNs, addressing the critical challenges of restricted gradient flow in deep, multi-layered QuNNs without using any additional qubits. We introduce trainable quanvolutional layers, significantly increasing the scalability and learning potential of QuNNs. To resolve complexities in gradient-based optimization for QuNNs with multiple layers, we propose ResQuNNs that utilize residual learning and skip connections to facilitate gradient flow across multiple trainable quanvolutional layers. While QResNets enhance frequency generation and flexibility in quantum circuits, our work specifically targets the scalability and learning potential of QuNNs, providing practical solutions for efficient training in deep quantum neural networks.
III ResQuNN Methodology
@@
III-A Trainable Quanvolutional Layers in QuNNs.
The first step is to make the quanvolutional layers trainable. The majority of works in the literature utilizes these layers for efficient feature extraction and pass them to classical layers, however, the quanvolution operation itself is not trainable. In order to enhance the adaptability of these layers within QuNNs it is important to make these layers trainable to harness their full potential. Making the quanvolutional layer trainable entails making the quantum circuit (labeled as 1 in Fig. 5) trainable. We do this by making the quanvolution layer as a pytorch layer which is later used as a standard layer during the model construction. The trainable QuNN utilized in this study is depicted in purple shaded region of Fig. 5, and operates through the following steps.
@@
    where $l$ represents the total number of layers, i.e., number of times the circuit is repeated before measurement. The size of output after every quanvloutional layers is detailed in Table II. |$\ket{\psi_{\text{QCL1}}}$ is the final state state after the first quanvolutionla layer before measurement and upon measurement the quantum state collapses and produce a classical result:
@@
    First, we use classical layer to postprocess the output of quanvolutional layer, as highlighted in label 2 in Fig. 5). This is a major practice in state-of-the-art QuNN architectures, where a fully connected neuron layer is used to take the output of the quanvolutional layer and produces the final classification output. The number of neurons in this layer are equal to the total number of classes in the dataset used, which in our case is $10$ since we used MNIST handwritten digit dataset which has $10$ classes (more details in Section IV).
    Let’s assume the output of the quanvolutional layer is a set of measurements $\{M=m_{1},m_{2}\ldots m_{n}\}$, where each $m_{i}$ represents the measurement of a qubit in terms of classical information, such as the expectation values in computational basis. This result is transformed to a vector format and then passed to the fully connected layer with $10$ neurons. The output feature size from quanvolutional layers is different and depends on the kernel size and stride similar to that of classical convolutional layers. We experiment with a kernel size of $2\times 2$ and stride of $2$. The output feature size after each quanvolutional layer in case of different residual configuration (label 5 in Fig. 5) is presented in the Table II.
    TABLE II: Output Feature Sizes for Two Quanvolutional Layer
@@
    The classical neural network layer with $10$ neurons is a fully connected layer, where each neuron receives inputs from all $n$ measurements of the quanvolutional layer and processes these inputs to produce an output. The output of $j$-th neuron in the classical layer can be described by the following Equation:
@@
    We also use quantum circuit to postprocess the result from quanvolutional layer, as highlighted by label 3 in Fig. 5. This is primarily for two reasons: firstly to demonstrate the working of fully quantum QuNNs, and secondly, to better understand the importance of residual connections by removing the influence of classical layers being used at the end.
    The use of a quantum circuit for post-processing introduces certain computational overheads, mainly due to the prerequisite of re-encoding the (classical) outputs from the quanvolutional layers, a process that entails considerable computational expense. Also the following trainable quantum circuit for postprocessing quanvolutional layer’s results after the encoding takes longer to produce the results than a simple classical neuron layer. We use $4$ classes from MNIST dataset for experimental simplification. For the purpose of encoding the outputs from the quanvolutional layer, Amplitude Encoding [26], is used which can be described by the following Equation:
@@
    where $x_{i}$ is the $i^{th}$ element of $x$ and |$\ket{i}$ is the $i^{th}$ computational basis state. We used amplitude encoding for postprocessing the results of quanvolutional layers primarily because it allows to encode $2^{n}$ features in $n$ qubits. Based on Table II, the smallest output feature size from quanvolutional layer(s) is $7\times 7\times 1=49$, which requires atleast $49$ qubits via angle encoding (one qubit per feature) and only $6$ qubits via amplitude encoding. Similarly, the maximum output size of quanvolutional layer is $28\times 28\times 1=784$, requiring $784$ qubits, if encoded in qubit rotation angles, whereas Amplitude Encoding reduces the requirement to merely $10$ qubits. The use of $784$ qubits is not practical in current NISQ devices, therefore we use amplitude encoding to encode the qunavolutional layer’s result for further processing by a quantum circuit. To ensure a consistent impact of the quantum circuit’s post-processing across all residual configurations, a $10$-qubit circuit is utilized for processing the outputs of the quanvolutional layer, corresponding to the maximum feature size of $784$. However, as mentioned earlier, we use $4$ classes, therefore, only $4$ qubits are measured each corresponding to one class. The output of quanvolutional, once encoded, is subjected to the unitary evolution through a simple quantum circuit, which can be described by the following Equation:
@@
III-B Proposed Residual Quanvolutional Neural Networks (ResQuNNs).
We observe that optimizing multiple (two or more) quanvolutional layers presents a significant challenges. One of the main difficulties lies in gradients accessibility across these layers during training, which are essential in gradient-driven optimization techniques. We typically notice that when we train two or more quanvolutional layers, the gradients of only the last layer are accessible. We have shown these results as a primary motivation of this paper in Section I-A.
To overcome the gradient accessibility issue encountered in multi-layered QuNNs, we introduce ResQuNNs. Drawing inspiration from classical residual neural networks, ResQuNNs incorporate residual blocks between quanvolutional layers. The residual blocks combine the output of existing quanvolutional layers with their own or the previous layer’s input before forwarding it to the subsequent quanvolutional layer, as shown in label 4 in Fig. 5. These residual connections play a pivotal role in facilitating comprehensive gradient access and, in turn, improving the overall training performance of QuNNs. The integration of residual blocks ensures that gradients flow efficiently across the network, thereby overcoming the challenges associated with training multiple quanvolutional layers.
III-C Strategic Placement of Residual Blocks.
Within the proposed ResQuNN architecture, not every random residual configuration results in comprehensive gradient flow throughout the network, and the residual blocks needs to be strategically placed between quanvolutional layers. In our empirical investigation, we conduct extensive experimentation to determine the appropriate locations for inserting residual blocks within ResQuNNs comprising two quanvolutional layers, as shown in Fig. 5, labeled 5. This analysis helps us uncover the most effective residual configurations for ResQuNNs, enabling the gradients to flow throughout the proposed ResQuNN architecture, resulting in a deeper architecture and eventually the enhanced training performance. Based on the results of two quanvolutional layers, only the most effective configurations, i.e., those facilitating gradient flow through both layers were chosen for subsequent examination within three-layer quanvolutional frameworks, as denoted by label 6 in Fig. 5.
We then trained ResQuNNs with two quanvolutional layers for all residual configurations for a multiclass classification problem (more on dataset in Section IV) to assess the impact of gradient accessibility, both throughout the entire network (all quanvolutional layers) and within segments of the network (only the last quanvolutional layer), on the learning efficacy of QuNNs.
TABLE III: Dataset and Hyperparemeters Specifications. BS = batch size, Opt = optimizer and LR = learning rate
@@
IV Experimental Setup
An overview of the experimental toolflow utilized in this paper is depicted in Fig. 6. We trained the proposed ResQuNNs on a selected subset of the MNIST dataset. Table III summarizes the dataset and hyperparameters specification used in this paper. The ResQuNNs are trained for $30$ epochs. Adam optimizer with learning rate of $0.01$ is used for the optimization. The objective function chosen for training is cross entropy loss, which can be described (for C classes) by the following equation: